OpenStudy (lilmike234):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4 or 11.25. What is the common ratio in this sequence?

OpenStudy (campbell_st):

well a term in a geometric sequence is found using \[a_{n} = a \times r^{n -1}\] and you know the 2nd term \[20 = a \times r^{2 -1} ~~~or~~~20 = a \times r\] and he 4th term \[\frac{45}{4} = a \times r^3\] which can be written as \[\frac{45}{4} = a \times r \times r^2\] make the substitution for the 2nd term \[\frac{45}{4} = 20 \times r^2\] now you can solve for r

OpenStudy (campbell_st):

and there are 2 possible answerz for r

OpenStudy (lilmike234):

Which are 1/7 and 3 right ?

OpenStudy (campbell_st):

 so then simplify the fraction